OpenStudy (anonymous):

College Algebra: reviewing for Final Exam! If f(x)=x^2+2x, then find f(a+h)-f(a). What's the first thing I do?

OpenStudy (jhannybean):

find f(a+h) by replacing your `x` with `a+h`

OpenStudy (jhannybean):

next find f(a) by replacing `x` with `a`

OpenStudy (anonymous):

So, it would look like this: \[f(a-h)=(a-h)^{2}+2(a-h)\]

OpenStudy (anonymous):

and \[f(a)=a^{2}+2a\]

OpenStudy (jhannybean):

it's not `f(a-h)` it's `f(a + h)`

OpenStudy (anonymous):

oops lolx

OpenStudy (anonymous):

\[f(a+h)=a ^{2}+h ^{2}+2a+2h\]

OpenStudy (anonymous):

Would I add f(a)=a^2+2a to that equation?

OpenStudy (jhannybean):

f( a + h) = (a+h)^2 +2(a+h) = \(a^2 +2ah+h^2 + 2a +2h\)

OpenStudy (jhannybean):

Do you see your mistake?

OpenStudy (anonymous):

So you did this: